OpenStudy (shaniehh):

If the polynomial P(x) has roots −5, −1, 2, and 2, which of the following represents the factored form of function P(x)?

OpenStudy (shaniehh):

A, (x − 5)(x − 1)(x + 2)(x + 2) B.(x + 5)(x + 1)(x − 2)(x − 2) C.(x − 5)(x − 1)(x + 2) D.(x + 5)(x + 1)(x − 2)

Directrix (directrix):

If these are roots: −5, −1, 2, and 2, then they showed up as x = -5, x = -1, x = 2, and x = 2 That came from: x + 5 = 0 , x + 1 =0, x-2 =0 and x -2 =0 which came from (x + 5)(x + 1)(x − 2)(x − 2) = 0

Directrix (directrix):

@shaniehh

OpenStudy (shaniehh):

thanks i got it aready

Directrix (directrix):

Okay.